MATRIX SPECTROSCOPY OF PEROXYL RADICALS VIA A HYPERTHERMAL NOZZLE

\underline{EVAN B. JOCHNOWITZ}, X. ZHANG, G. B. ELLISON, Dept. Chemistry, University of Colorado Boulder, CO 80309-0215; M. R. NIMLOS, Center for Renewable Chemical Technologies & Materials, NREL, 1617 Cole Blvd., Golden, CO 80401.

We have developed a hyperthermal nozzle to produce organic radicals for study in a cryogenic matrix. This hot nozzle thermally decomposes organic precursors in a stream of Ar and produces roughly 10^{13} hydrocarbon radicals/pulse. This technique has enabled us to prepare alkylperoxyl radicals by co-deposition with oxygen. In the atmospheric ?processing? of organic aerosols, surface bound peroxyl radicals will be important reaction intermediates. We have successfully studied the CH3OO radical by combining CH_{3} and O_{2} in an Ar matrix at 20 K. Preliminary results for allylperoxyl (CH_{2}CHCH_{2}OO) and propargylperoxyl (HCCCH_{2}OO) radical will be presented.